An office supply company manufactures and sells X permanent markers per year at a price of P €/unit. The Price/Demand equation
for the markers is:
1
Expert's answer
2020-06-23T19:48:31-0400

"P=7-0.002x"


"TC=1000+2x"

"TR=PQ" (Total revenue= Price "*" Quantity

"(7-0.002x" ")x"

"MR=" "dTR\over dQ" "=7x-0.002x{^2}"

"=7-0.004x"

"MC=" "dTC\over dQ" "=1000+2x"

"=2"

For profit maximization, marginal revenue = marginal cost.

"MC=MR"

"2=7-0.004x"

"x=1250"

"Price"

"P=7-0.002(1250)"

"=4.50"